OpenStudy (anonymous):

Three brothers spent different amounts of money on their recent vacation. Matt spent $25 less than Todd. Jeff spent 3 times as much as Matt. How much did each brother spend if they spent a total of $275?

jimthompson5910 (jim_thompson5910):

Let M = amount of money Matt spent J = amount of money Jeff spent T = amount of money Todd spent

jimthompson5910 (jim_thompson5910):

if "Matt spent $25 less than Todd", then we cay say that M = T - 25

jimthompson5910 (jim_thompson5910):

if "Jeff spent 3 times as much as Matt", then we know J = 3*M

jimthompson5910 (jim_thompson5910):

"they spent a total of $275", so we can say M+J +T = 275

OpenStudy (anonymous):

how do we find the variables?

jimthompson5910 (jim_thompson5910):

now use the 3 equations together to solve M+J+T = 275 ... start with the 3rd equation M+3M+T = 275 ... plug in J = 3M (replace J with 3M) 4M+T = 275 ... Combine like terms. 4(T-25)+T = 275 ... plug in M = T - 25 (replace M with T-25) 4T-100+T = 275 ... distribute I'll let you finish up

jimthompson5910 (jim_thompson5910):

notice how all 3 equations were involved and how we went from an equation with 3 variables to an equation with one variable

OpenStudy (anonymous):

so does todd equal x?

jimthompson5910 (jim_thompson5910):

todd is T

jimthompson5910 (jim_thompson5910):

what do you get when you solve 4T-100+T = 275 for T

OpenStudy (anonymous):

1275

jimthompson5910 (jim_thompson5910):

that's too high

OpenStudy (anonymous):

5t-100=275

jimthompson5910 (jim_thompson5910):

so far, so good, keep going

OpenStudy (anonymous):

its a decimal right?

jimthompson5910 (jim_thompson5910):

no, it's a whole number

OpenStudy (anonymous):

i got 25.4..

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

sorry i typed 27 lol

jimthompson5910 (jim_thompson5910):

5t-100=275 5t-100+100=275+100 ... Add 100 to both sides. 5t+0 = 375 ... Combine like terms. 5t = 375 t = ???

jimthompson5910 (jim_thompson5910):

ah i gotcha

OpenStudy (anonymous):

t=75.

jimthompson5910 (jim_thompson5910):

good, so Todd spent $75

jimthompson5910 (jim_thompson5910):

now that you know that T = 75, you can use this and the other equations to find M and J

OpenStudy (anonymous):

so we add all the answers up or subtract?

jimthompson5910 (jim_thompson5910):

what were the 3 original equations

jimthompson5910 (jim_thompson5910):

since T = 75 and we have the equation M = T - 25, we can say this M = T - 25 M = 75 - 25 .... replace T with 75 (since T = 75) M = 50 So Matt spent $50

OpenStudy (anonymous):

m= t-25 3xm m+j+t

OpenStudy (anonymous):

but i thought matt was 45?

jimthompson5910 (jim_thompson5910):

J = 3*M J = 3*50 ... replace M with 50 (since we now know that M = 50) J = 150 So Jeff spent $150

jimthompson5910 (jim_thompson5910):

Summary: Todd spent $75 Matt spent $50 Jeff spent $150

jimthompson5910 (jim_thompson5910):

not sure how you got $45

OpenStudy (anonymous):

damn girllllll..... you shouldnt be talking to this loser you should kik me ;)

OpenStudy (anonymous):

i thought matt's equation was m-25,

jimthompson5910 (jim_thompson5910):

SwoleNerd this is a study site, not facebook

OpenStudy (anonymous):

and i dont have kik.

OpenStudy (anonymous):

lol girl how old are you im 15

OpenStudy (anonymous):

ok jim i have another question.

jimthompson5910 (jim_thompson5910):

Matt's equation is M = T - 25

jimthompson5910 (jim_thompson5910):

replace T with 75 (remember we found that T = 75) and you'll get M = 50

OpenStudy (anonymous):

ok i got it.

OpenStudy (anonymous):

god bless!

jimthompson5910 (jim_thompson5910):

what's your other question?

OpenStudy (anonymous):

its really long so hold on.

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

At the end of the lacrosse season, the coach always posts the names of the top 5 scorers and the number of goals each person scored throughout the season. this year, the coach, who is also a math teacher, decided to have some fun with his players. this is what he posted.

OpenStudy (anonymous):

Erin scored the least number of goals. Macy scored 4 times as many goals as Erin. Carly scored 2 goals less than Macy. Shelly scored one half as many goals as carly. Molly scored 1 more goal than Macy.

OpenStudy (anonymous):

If the total number of team goals for the season was 253, determine how many goals each of the top 5 players scored.

OpenStudy (anonymous):

hello

jimthompson5910 (jim_thompson5910):

sorry still reading and thinking

jimthompson5910 (jim_thompson5910):

let's start by setting up variables Let E = number of goals Erin scored Ma = number of goals Macy scored C = number of goals Carly scored S = number of goals Shelly scored Mo = number of goals Molly scored

jimthompson5910 (jim_thompson5910):

"Macy scored 4 times as many goals as Erin." means Ma = 4*E

jimthompson5910 (jim_thompson5910):

"Carly scored 2 goals less than Macy." means C = Ma - 2

jimthompson5910 (jim_thompson5910):

"Shelly scored one half as many goals as carly." means S = C/2

jimthompson5910 (jim_thompson5910):

"Molly scored 1 more goal than Macy." means Mo = Ma + 1

jimthompson5910 (jim_thompson5910):

"the total number of team goals for the season was 253" so we know that E + Ma + C + S + Mo = 253

jimthompson5910 (jim_thompson5910):

because Ma = 4*E, we can solve for E to get E = Ma/4

jimthompson5910 (jim_thompson5910):

E + Ma + C + S + Mo = 253 E + Ma + C + S + Ma + 1 = 253 ... plug in Mo = Ma + 1 E + Ma + C + C/2 + Ma + 1 = 253... plug in S = C/2 E + Ma + Ma - 2 + (Ma - 2)/2 + Ma + 1 = 253... plug in C = Ma - 2 Ma/4 + Ma + Ma - 2 + (Ma - 2)/2 + Ma + 1 = 253... plug in E = Ma/4 Ma/4 + 3*Ma + (Ma - 2)/2 - 1 = 253 Ma/4 + 3*Ma + Ma/2 - 2/2 - 1 = 253 Ma/4 + 3*Ma + Ma/2 - 1 - 1 = 253 Ma/4 + 3*Ma + Ma/2 - 2 = 253 Ma/4 + 12*Ma/4 + 2Ma/4 - 2 = 253 (Ma + 12*Ma + 2Ma)/4 - 2 = 253 (15Ma)/4 - 2 = 253 (15Ma)/4 = 253+2 (15Ma)/4 = 255 15Ma = 255*4 15Ma = 1020 Ma = 1020/15 Ma = 68 So Macy scored 68 points E = Ma/4 E = 68/4 E = 17 Erin scored 17 points Mo = Ma + 1 Mo = 68 + 1 Mo = 69 Molly scored 69 points C = Ma - 2 C = 68 - 2 C = 66 Carly scored 66 points S = C/2 S = 66/2 S = 33 Shelly scored 33 points Summary: E = 17 Ma = 68 C = 66 S = 33 Mo = 69